Substitution method:
3x-6y=-2
2x-3y=1

Answer by josgarithmetic(39618)

Pick an equation and solve for either variable.
3x-6y=-2
3x=6y-2
x=2y-2/3
Substitute into the other equation.
2(x)-3y=1
2(2y-2/3)-3y=1
4y-4/3-3y=1
y-4/3=1
y=4/3+1
y=7/3 and use this to find the value for x.